A toy rocket shot vertically upward with an initial velocity of 160 feet per second (fps) has a time t a velocity given by the equation v = 160-32t, where air resistance is neglected. In how many seconds will the rocket reach its highest point? (Hint: Find t when v = O)
Answer by fcabanski(1391)

v = 160-32t = 0


160 = 32t


32t=160


t = 160/32 = 5 seconds